In 8 hours, two campers rowed 15 miles down a river and then rowed back to their campsite. The rate of the river's current was 1 mph. Find the rate at which the campers row in calm water.
What I know is that it was 8 hours total and it took 15 mi there and back.
Please help! I've been trying to solve it for now an hour and 31 minutes.

Answer by mananth(16946)

let the speed of the boat be xmph in still water
down stream the speed will be x+1 mph
time taken down stream = 15 / x+1 hours
upstream the speed will be x-1 mph
time taken upstream = 15 / x-1
sum of these two times = 8 hours
15 / x+1 + 15 /x-1 = 8
15(x-1)+15(x+1) / x^2-1= 8
15x-15 +15x +15 = 8(x^2-1)
30x = 8x^2-8
8x^2-30x-8 =0
8x^2-32x+2x-8=0
8x(x-4)+2(x-4)=0
(x-4)(8x+2)=0
x= 4 or -2/8
the rowing speed in still waters = 4 mph